OpenStudy (anonymous):

CHECKING MY ANSWER: a company uses two vans to transport workers from a free parking lot to the workplace between 7:00 and 9:00a.m. one van has 6 more seats than the other. the smaller van makes two trips every morning while the larger one makes only one trip. the two vans can transport 57 people, maximum. Let x be the seats in the small van while y be the seats for the larger van. how many seats does the larger van have?

OpenStudy (anonymous):

23 seats?

OpenStudy (anonymous):

@Preetha @Michele_Laino

OpenStudy (anonymous):

@amistre64 (:

OpenStudy (stormswan):

how did you figure this out?

OpenStudy (anonymous):

x-6+x-6+x=57 x=23

OpenStudy (stormswan):

I got 15 from your equation

OpenStudy (amistre64):

small + large = 57 if 23 is the large, 23+23 = 46, 2 large vans cant even transport 57 then

OpenStudy (stormswan):

he was saying that the small was 23.

OpenStudy (amistre64):

still doesnt work for me 23+23+6 is not 57

OpenStudy (amistre64):

large = small + 6 small + small + 6 has to be 57

OpenStudy (stormswan):

but wouldn't it be 23+23+29? Because the small van takes 2 trips and then the large is +6

OpenStudy (anonymous):

Yesh..?

OpenStudy (amistre64):

2 trips ... too many words to keep track of lol

OpenStudy (anonymous):

So I'm right?

OpenStudy (amistre64):

:) one van has 6 more seats than the other. large = small + 6 the smaller van makes two trips every morning small + small while the larger one makes only one trip. small + 6 the two vans can transport 57 people 3small +6 = 57

OpenStudy (amistre64):

can we assume large has more seats than small?

OpenStudy (anonymous):

Yea, 6 more,.

OpenStudy (amistre64):

ive seen a lot of clowns come out of a vw ... just saying :)

OpenStudy (anonymous):

xDD

OpenStudy (amistre64):

51/3 = 17 for the small, so 6 more is 23 is fine to mee

OpenStudy (anonymous):

So the question asks for the larger van, It holds 29 right

OpenStudy (stormswan):

small= 17 large= 23

OpenStudy (amistre64):

of course not. small is 17, 17+6 is 23

OpenStudy (anonymous):

Oh yeah, Gotcha.